# Towards autonomous photogrammetric forest inventory using a lightweight under-canopy robotic drone

### v1

@@ 2.2 DBH estimation from under-canopy drone data
Table 1: Accuracy of the stem diameter breast height (DBH) estimations in other studies. SfM-MVS: Structure-from-motion and multi-view stereo; ULS: Under canopy laser scanning; ALS: Above canopy laser scanning. *averaged values from three plots
Dataset Type Number of trees RMSE (cm) RMSEBias (cm) Bias%
Shimabuku et al. (2023) plot I Drone SfM-MVS 7 0.4 3.2 - -
Shimabuku et al. (2023) plot II Drone SfM-MVS 18 0.7 3.6 - -
@@
Liang et al. (2024) traj. I ULS 143 9.43 63.41 7.17 48.21
Liang et al. (2024) traj. II ULS 143 4.48 21.37 5.01 14.91
Hyyppä et al. (2020a) plot I ULS 42 0.6 2.3 0.34 1.3
Hyyppä et al. (2020a) plot II ULS 43 1.1 3.5 0.12 0.4
Wang et al. (2021) ULS&ALS - 7.95 33.35 3.72 15.61
Muhojoki et al. (2024) * ULS 50 1.50-1.60 6.70-6.90 0.10-0.70 0.60-3.00
@@ 3.1 Autonomous navigation algorithms
In this study, the objective was to develop a prototype drone capable of autonomously avoiding obstacles under the forest canopy and operating without GNSS. The drone was required to autonomously fly from the takeoff location to a specified goal point while recording sensor data during the flight. This prototype will later serve as a development platform for more extended and sophisticated data collection tasks.
Among the open-source solutions presented in Section 2.1, the solution by Zhou et al. (2022) was chosen as the base of the prototype. The solution was proven to work with only one stereo camera making it possible to build a smaller and lighter platform. Furthermore, cameras are typically lighter and more power-efficient than LiDARs (Reddy Cenkeramaddi et al. 2020) making them an attractive solution for small under-canopy drones, for which the weight and battery consumption must be optimized. This enables building a compact system that is capable of operating in dense forests.
The main details of the algorithms used in the chosen solution are described briefly in the following and the complete system architecture of the prototype is presented in Figure 1. For more detailed information the readers are referred to the cited original sources.
@@ 3.1.1 EGO-Planner-v2
The source code of the planner algorithm by Zhou et al. (2022) was named in Github as EGO-Planner-v2 (ZJU FAST Lab Team 2022a). Even though the EGO-Planner-v2 is designed for swarm flying, in this study, only a single drone system was used. The planner is running on top of the Robot Operating System (ROS) (Quigley et al. 2009). The mapping of EGO-Planner-v2 is based on probabilistic occupancy grid maps (Moravec and Elfes 1985). The mapping module uses depth images as an input and the system adopts the fixed-sized circular buffers proposed by Usenko et al. (2017) for maintaining the local map. The maps also have a virtual floor and a virtual ceiling defining the minimum and maximum altitudes of the grid map respectively. Virtual floor and virtual ceiling restrict the altitude where the algorithm is allowed to plan paths.
@@
where where $J_{x}$ are the penalty terms and $\lambda_{x}$ are the corresponding weights. Subscripts $x=\{s,t,d,o,u\}$ stand for the trajectory smoothness, total flying time, dynamical feasibility, obstacle avoidance, and the uniform distribution of constraint evaluation points along the trajectory, respectively. In addition to the above mentioned penalty terms, EGO-Planner-v2 includes also penalty terms for the swarm operations, but those are omitted here. A* is utilized as the global path-searching algorithm given as an input for the optimization algorithm.
@@ 3.1.2 VINS-Fusion
The VIO algorithm used by Zhou et al. (2022) was VINS-Fusion (Qin, Li, and Shen 2018; Qin et al. 2019b)(VINS stands for ”Visual-Inertial system”), which used grayscale stereo images and IMU data to track the position and orientation of the drone.
VINS-Fusion detects Shi-Tomasi corner features (Shi and Tomasi 1994), and the features are tracked between frames with Kanade–Lucas–Tomasi (KLT) Tracker (Lucas and Kanade 1981). IMU data is used for preintegration, i.e. position, velocity, and orientation are integrated from IMU measurements between two frames. The state optimization is performed within a sliding window including the position, velocity, rotation, and IMU biases. To reduce the computational complexity, old measurements are marginalized into a prior of the estimation process. Certain frames are also selected as keyframes which are saved to a global pose graph after being marginalized out from the sliding window. Keyframes are selected based on the number of tracked features and the average parallax between the features in the newest frame and the latest keyframe.
VINS-Fusion supports loop detection-based relocalization and global pose graph optimization to reduce the accumulated drift in the pose estimate. The loops are detected by storing additional corner features as BRIEF descriptors (Calonder et al. 2010) to DBoW2 (Galvez-López and Tardos 2012) bag-of-word database. After a loop detection, the relocalization process aligns the latest window to past poses and optimizes it utilizing also the feature correspondences with past poses. In addition to sliding window relocalization, also the global pose graph is optimized after loop detections to enforce the global consistency of the poses. Since the roll and pitch angle estimates are drift accumulation-free due to full observability from the gravity measurement of the IMU, the pose graph optimization is performed in 4 degrees of freedom.
VINS-Fusion also adopts an online temporal calibration method for inertial and visual measurements (Qin and Shen 2018). The method can be used for adjusting the time offset $t_{d}$ between the camera and the IMU online.
In the swarm operations of the original article (Zhou et al. 2022), the drones used also relative distances measured with Ultra Wide Band (UWB) sensors to correct the drift in the pose estimate of VINS-Fusion. However, with a single drone of this study that function is omitted.
The parameters used in the configuration of VINS-Fusion were chosen so that the real-time performance was guaranteed. In the initial tests, the IMU noise values used during the calibration were over-optimistic for VINS-Fusion causing instability to the estimates, so the IMU noise values were increased more. All VINS-Fusion configuration parameters are listed in Table 2.
@@ 3.1.3 Tracking controller
The original EGO-Planner-v2 article (Zhou et al. 2022) did not specify the utilized method for trajectory tracking. In this study, an open-source tracking controller (ZJU FAST Lab Team 2022b) by Zhejiang University was used. The tracking controller gets setpoints for position, velocity, acceleration, and yaw from the trajectory planner, and converts those to attitude and thrust setpoint. The setpoints are forwarded to PX4 flight controller software (PX4 Community 2023) responsible for low level attitude and angular rate controls.
@@ 3.2 Drone hardware
In the physical robot drone prototype, an onboard computer NVIDIA Jetson Orin NX (Santa Clara, California, USA) was used. The used PX4-compliant autopilot was Holybro Kakute H7 (Hongkong, China), which contained also the IMU. The navigation algorithms mentioned in the previous subsection ran on top of the ROS, and the Mavros-protocol was used for communication between the onboard computer and the autopilot. As a camera, Intel RealSense D435 (Intel Realsense Development Team 2023)(Keselman et al. 2017) was used. The camera was capturing both depth images for 3D mapping and grayscale stereo images for VIO.
@@
RealSense D435 is based on active stereoscopy and the camera has an infrared laser projector emitting dots that can be used for triangulation by stereo camera. This can improve the triangulation in low-light conditions and to homogeneous surfaces that contain no features. However, the dots of the laser emitter cause problems for VIO since they are moving with the camera. In the previous setup described by (Karjalainen et al. 2023), a modified version of ROS-RealSense driver was used by Zhou et al. (2021b) that turns the laser emitter of the camera constantly on and off making it possible to produce high-quality depth images in every other frame, and stereo images without the laser dots in every other frame. However, this solution comes with the drawback of halving the output frame rate of the camera.
@@
In the first test flights described by (Karjalainen et al. 2023), difficulties were encountered in detecting small branches with a depth camera resolution of $640\times 480$, leading to the decision to increase the depth camera resolution to a maximum value of $1280\times 720$. Stereo images for VIO were downsampled to resolution of $640\times 360$ to ensure real-time performance. As the maximum frame rate of D435 at a resolution of $1280\times 720$ is 30 Hz, the modified driver version would have resulted in a low frame rate of 15 Hz. To ensure the VIO performance in quick turns, the frame rate of 30 Hz was preferred, and hence disabled the IR emitter making the camera function as a passive IR stereo camera. The choice can be further justified with the fact that the usefulness of the IR emitter in forest conditions is questionable, since sunlight contains infrared light that can interfere with the emitted artificial features(Kuan, Ee, and Wei 2019), and the forests typically do not contain large homogeneous surfaces. The Realsense D435 onboard camera was configured to use the builtin automatic exposure control.
@@
The camera-IMU system was calibrated with an open-source camera calibration toolbox, kalibr (Furgale, Barfoot, and Sibley 2012)(Furgale, Rehder, and Siegwart 2013)(Maye, Furgale, and Siegwart 2013). In the calibration, the intrinsic and extrinsic parameters of the left and right grayscale cameras and the time shift between the camera and the autopilot IMU were calculated from recorded data where Aprilgrid (Olson 2011) was used as a calibration target. The IMU noise parameters needed for the calibration were first obtained with open source tool Allan Variance ROS (Oxford Dynamic Robot Systems Group 2022), and the obtained values were multiplied with 10 before the calibration as suggested in the project wiki of kalibr (ETH Zürich Autonomous Systems Lab 2023).
@@
The computer, the camera, and the autopilot were attached to a 330 mm drone frame. The propellers had two pieces of 6 cm long blades each. The measured weight of the drone including the motors and the onboard computer was 791 g without the batteries, and 1153 g with the batteries. An overview of the used drone hardware is presented in Figure 2.
@@ 3.3 Test sites
The performance of the system was evaluated in three different Boreal forest areas. The flight tests of the system were performed in two test areas located in Evo, Finland (61.19°N, 25.11°E). The flight paths went through well-documented forest sample plots that have been widely used in previous studies, such as in Liang et al. (2018, 2019); Wang et al. (2019); Hyyppä et al. (2020a, c), and are classified into three complexity categories, i.e., “easy”, “medium”, and “difficult” based on, mainly tree density and amount of the understory vegetation in the plots.
The first plot has been classified as ”medium”, and the second one has been classified as ”difficult”. The medium difficulty plot had a tree density of 650 trees/ha and a mean diameter at breast height (DBH) of 28 cm. The most common tree species in the plot were Norway spruce (Picea abies (L.) H.Karst.)(81.2%) and Scots pine (Pinus sylvestris L.)(9.4%). The difficult forest had a tree density of 2000 trees/ha and a mean DBH of 17 cm. The most common tree species in the plot were Norway spruce (64.4%) and aspen (Populus tremula, L.) (18.3%). Only trees with a DBH greater than 5 cm were included in these statistics. From now on, these two test areas are called ”Evo-medium” and ”Evo-difficult”, respectively.
For georeferencing purposes, circular coded targets provided by Agisoft Meatahsape (Agisoft LLC, St. Petersburg, Russia) were placed within the forest plots. The targets had a diameter of 24.7 cm, and the middle dots had a size of 7 cm. Relative distances between the coded targets were measured with Leica Nova TS60 total station with a precision of approximately 2mm (Leica Geosystems 2020). Figure 5 shows coded targets in ”Evo-difficult” forest during the test flights.
The third test area was a spruce forest in Paloheinä, Finland (60.26°N, 24.92°E). Aim of the experiment there was to test the loop detection performance of VINS-Fusion by walking with the drone in an environment without coded targets. Estimated average tree density in the area was approximately 2000 trees/ha.
Figure 3: View from the first takeoff location in the Evo-medium test area. The takeoff location was changed after five test flights.
Figure 4: View from the takeoff location in the Evo-difficult test forest

### v2

@@ 2.1 DBH estimation from under-canopy drone data
Table 1: Accuracy of the stem diameter breast height (DBH) estimations in other studies. SfM-MVS: Structure-from-motion and multi-view stereo; ULS: Under canopy laser scanning; ALS: Above canopy laser scanning. *averaged values from three plots. The study by He et al. (2025) was omitted from this table since they reported only median errors instead of RMSE.
Dataset Type Number of trees RMSE (cm) RMSE % Bias (cm) Bias %
Chisholm et al. (2013) ULS 12 - 25.1 - -
Liang et al. (2024) traj. I ULS 143 9.79 61.41 7.18 45.03
Liang et al. (2024) traj. II ULS 143 5.13 22.01 3.40 14.62
Hyyppä et al. (2020a) plot I ULS 42 0.69 2.2 0.31 1.1
Hyyppä et al. (2020a) plot II ULS 43 0.92 3.1 0.29 1.0
Hyyppä et al. (2020c) plot I ULS 42 0.6 2.3 0.34 1.3
Hyyppä et al. (2020c) plot II ULS 43 1.1 3.5 0.12 0.4
Wang et al. (2021) ULS&ALS - 7.95 33.35 3.72 15.61
Muhojoki et al. (2024) * ULS 50 1.50-1.60 6.70-6.90 0.10-0.70 0.60-3.00
Prabhu et al. (2024) ULS 20 1.67 7.42 - -
Cheng et al. (2024) Plot I ULS 77 3.8 12.8 - -
Cheng et al. (2024) Plot II ULS 20 2.0 7.5 - -
Shimabuku et al. (2023) plot I Drone SfM-MVS 18 0.7 3.6 - -
Shimabuku et al. (2023) plot II Drone SfM-MVS 7 0.4 3.2 - -
@@
The highest estimation accuracies have been obtained with large systems unable to fly in dense forests or with test plots containing very few reference trees, but miniaturized systems have still yielded promising results. Notably, all studies utilizing SfM point clouds have been performed in environments that differ significantly from boreal forests.
2.2 Autonomous flying inside forests
Even though only a few studies have yet utilized robotic under-canopy drones in DBH estimation, drone autonomy itself has been an actively researched topic in the robotics community during the past few years. Recent advances in mobile robotics and the enhanced computational capabilities of small embedded computers have made it realistic to build drones capable of real-time trajectory planning in GNSS-denied cluttered environments. Although technology is still in the early stage of development, some recent studies have proposed various methods that have been successfully tested in forest environments. Most importantly, the authors of these methods have also shared their source code to be openly available for the research community to encourage more development in the field. While the methods utilize various sensor setups, they can be divided into LiDAR-based methods (Liu et al. 2022; Prabhu et al. 2024; Ren et al. 2022; Liu, Ren, and Zhang 2024; Ahmad et al. 2022) and camera-based methods (Campos-Macías et al. 2021; Zhou et al. 2022; Loquercio et al. 2021; Nguyen et al. 2023) based on the sensor utilized for the sensing of the obstacles in the flight path. Since GNSS is not a suitable method for localization under the forest canopy, these methods rely on various Simultaneous localization and mapping (SLAM) or odometry algorithms and onboard sensor data to obtain the pose estimate in a local coordinate system. Most of these methods tackle the problem of collision-free trajectory planning using 3D grid maps, but the solutions by Loquercio et al. (2021) and Nguyen et al. (2023) replaced individual mapping and trajectory planning subtasks by deep neural networks mapping the depth images directly to collision-free trajectories. In addition, the methods proposed by Zhou et al. (2022) and Ahmad et al. (2022) concentrated a swarm of drones instead of autonomous flying of an individual drone.
However, most of these methods are typically validated by a single successful autonomous test flight in a specific forest environment. Only four of the articles reported multiple test flights inside real forests (Loquercio et al. 2021; Ren et al. 2022; Liu, Ren, and Zhang 2024; Prabhu et al. 2024). From these, (Loquercio et al. 2021) and (Prabhu et al. 2024) were the only methods tested at least in two different forest types. Moreover, since methods by (Loquercio et al. 2021), (Ren et al. 2022), and (Liu, Ren, and Zhang 2024) were all targeting high-speed flight, the test forests reported in those studies were relatively sparse, with a low amount of low-hanging branches and understory vegetation. The autonomous trajectory planning method by Prabhu et al. (2024) was also tested only in environments where the space between the trees was larger than in typical forests, and the data collecting flights in denser forests were manually piloted.
Therefore, open questions remain regarding the applicability and reliability of these algorithms in diverse forest environments, especially in challenging boreal forests.
@@ 3.1 Autonomous navigation algorithms
In this study, the objective was to develop a prototype lightweight camera-based drone capable of autonomously avoiding obstacles under the forest canopy and operating without GNSS. The drone was required to fly autonomously from the takeoff location to a specified goal point while recording stereo-camera data during the flight. This prototype will later serve as a development platform for more extended and sophisticated data collection tasks.
Among the open-source solutions presented in Section 2.2, the solution by Zhou et al. (2022) was chosen as the base of the prototype. The solution was the only open-sourced camera-based solution proven to work with only one stereo camera making it possible to build a smaller and lighter platform. Additionally, the test environment reported in the original article (dense bamboo forest) was the most difficult among the proposed solutions, even though it had a highly different structure than the typical boreal forests.
The main details of the algorithms used in the chosen solution are briefly described in the following, and the complete system architecture of the prototype is presented in Figure 1. For more detailed information, the readers are referred to the cited original sources.
@@ 3.1.1 EGO-Planner-v2
The source code of the planner algorithm by Zhou et al. (2022) was named in Github as EGO-Planner-v2 (ZJU FAST Lab Team 2022a). Although EGO-Planner-v2 is designed for swarm flying, in this study, the algorithm was modified for a single drone system to focus on navigation performance in challenging forest environments. The planner runs on top of the Robot Operating System (ROS) (Quigley et al. 2009). The mapping of EGO-Planner-v2 is based on probabilistic occupancy grid maps (Moravec and Elfes 1985). The mapping module uses depth images as an input, and the system adopts the fixed-sized circular buffers proposed by Usenko et al. (2017) for maintaining the local map. The maps also have a virtual floor and a virtual ceiling defining the minimum and maximum altitudes of the grid map, respectively. The virtual floor and the virtual ceiling restrict the altitude at which the algorithm can plan the paths.
@@
where $J_{x}$ are the penalty terms and $\lambda_{x}$ are the corresponding weights. Subscripts $x=\{s,t,d,o,u\}$ stand for the trajectory smoothness, total flying time, dynamical feasibility, obstacle avoidance, and the uniform distribution of constraint evaluation points along the trajectory, respectively. In addition to the above-mentioned penalty terms, EGO-Planner-v2 also includes penalty terms for the swarm operations, but those are omitted here. A* is utilized as the global path-search algorithm given as input for the optimization algorithm.
@@ 3.1.2 VINS-Fusion
The Visual-inertial odometry (VIO) algorithm used by Zhou et al. (2022) was VINS-Fusion (Qin, Li, and Shen 2018; Qin et al. 2019b)(VINS stands for ’Visual-Inertial system’), which used grayscale stereo images and IMU data to track the position and orientation of the drone.
VINS-Fusion detects Shi-Tomasi corner features (Shi and Tomasi 1994), and the features are tracked between frames with Kanade–Lucas–Tomasi (KLT) Tracker (Lucas and Kanade 1981). IMU data is used for preintegration, i.e., position, velocity, and orientation are integrated from IMU measurements between two frames. The state optimization is performed within a sliding window, including the position, velocity, rotation, and IMU biases. To reduce the computational complexity, old measurements are marginalized into a prior of the estimation process. Certain frames are also selected as keyframes, which are saved to a global pose graph after being marginalized out from the sliding window. Keyframes are selected by considering the number of tracked features and the average parallax between these features in the most recent camera frame and the latest keyframe.
VINS-Fusion supports loop detection-based relocalization and global pose graph optimization to reduce the accumulated drift in the pose estimate. The loops are detected by storing additional corner features as BRIEF descriptors (Calonder et al. 2010) to DBoW2 (Galvez-López and Tardos 2012) bag-of-word database. After a loop detection, the relocalization process aligns the latest window to past poses and optimizes it utilizing also the feature correspondences with past poses. In addition to sliding window relocalization, the global pose graph is also optimized after loop detections to enforce the global consistency of the poses. Since the roll and pitch angle estimates are drift accumulation-free due to full observability from the gravity measurement of the IMU, the pose graph optimization is performed in 4 degrees of freedom.
VINS-Fusion also adopts an online temporal calibration method for inertial and visual measurements (Qin and Shen 2018). The method can be used to adjust the time offset $t_{d}$ between the camera and the IMU online.
In the swarm operations of the original article (Zhou et al. 2022), the drones used also relative distances measured with Ultra Wide Band (UWB) sensors to correct the drift in the pose estimate of VINS-Fusion. However, with a single drone system in this study, that function is omitted.
The parameters used in the configuration of VINS-Fusion were chosen so that real-time performance was guaranteed. In the initial tests, the IMU noise values used during the calibration were over-optimistic for VINS-Fusion causing instability to the estimates, so the IMU noise values were increased more. All VINS-Fusion configuration parameters are listed in Table 2.
@@ 3.1.3 Tracking controller
The original EGO-Planner-v2 article (Zhou et al. 2022) did not specify the utilized method for trajectory tracking. In this study, an open-source tracking controller (ZJU FAST Lab Team 2022b) by Zhejiang University was used. The tracking controller gets setpoints for position, velocity, acceleration, and yaw from the trajectory planner and converts those to attitude and thrust setpoints. The setpoints are forwarded to the PX4 flight controller software (PX4 Community 2023) responsible for the low-level attitude and angular rate controls.
@@ 3.2 Drone hardware
In the physical robot drone prototype, an onboard computer NVIDIA Jetson Orin NX (Santa Clara, California, USA) was used. The used PX4-compliant autopilot was Holybro Kakute H7 (Hongkong, China), which also contained the IMU. The navigation algorithms mentioned in the previous subsection ran on top of the ROS, and the Mavros protocol was used for communication between the onboard computer and the autopilot. As a camera, Intel RealSense D435 (Intel Realsense Development Team 2023)(Keselman et al. 2017) was used. The camera was capturing both depth images for 3D mapping and grayscale stereo images for VIO.
@@
RealSense D435 is based on active stereoscopy, and the camera has an infrared laser projector emitting dots that can be used for triangulation by the stereo camera. This can improve the triangulation in low-light conditions and to homogeneous surfaces that contain no features. However, the dots of the laser emitter cause problems for VIO since they are moving with the camera. In the previous set-up described by (Karjalainen et al. 2023), a modified version of the ROS-RealSense driver by Zhou et al. (2021b) was used. The modified driver turns the laser emitter of the camera constantly on and off, making it possible to produce high-quality depth images in every other frame and stereo images without the laser dots in every other frame. However, this solution comes with the drawback of halving the output frame rate of the camera.
@@
In the first test flights described by (Karjalainen et al. 2023), difficulties were encountered in detecting small branches with a depth camera resolution of $640\times 480$, leading to the decision to increase the depth camera resolution to a maximum value of $1280\times 720$. The stereo images for VIO were downsampled at a resolution of $640\times 360$ to ensure real-time performance. As the maximum frame rate of D435 at a resolution of $1280\times 720$ is 30 Hz, the modified driver version would have resulted in a low frame rate of 15 Hz. To ensure the VIO performance in quick turns, the frame rate of 30 Hz was preferred, and hence, the IR emitter was disabled, making the camera function as a passive IR stereo camera. The choice can be further justified with the fact that the usefulness of the IR emitter in forest conditions is questionable since sunlight contains infrared light that can interfere with the emitted artificial features(Kuan, Ee, and Wei 2019), and the forests typically do not contain large homogeneous surfaces. The Realsense D435 onboard camera was configured to use the built-in automatic exposure control.
@@
The camera-IMU system was calibrated with an open-source camera calibration toolbox, kalibr (Furgale, Barfoot, and Sibley 2012; Furgale, Rehder, and Siegwart 2013; Maye, Furgale, and Siegwart 2013). In the calibration, the intrinsic and extrinsic parameters of the left and right grayscale cameras and the time shift between the camera and the autopilot IMU were calculated from recorded data where Aprilgrid (Olson 2011) was used as a calibration target. The IMU noise parameters needed for the calibration were first obtained with the open source tool Allan Variance ROS (Oxford Dynamic Robot Systems Group 2022), and the values obtained were multiplied by 10 before the calibration as suggested in the wiki of the kalibr project (ETH Zürich Autonomous Systems Lab 2023).
@@
The computer, the camera, and the autopilot were attached to a 330 mm drone frame. The propellers had two pieces of 6 cm long blades each. The measured weight of the drone, including the motors and the onboard computer, was 791 g without batteries and 1153 g with the batteries. An overview of drone hardware is presented in Figure 2.
@@ 3.3 Test sites
The performance of the system was evaluated in three different boreal forest areas. The flight tests of the system were performed in two test areas located in Evo, Finland (61.19°N, 25.11°E). The flight paths went through well-documented forest sample plots that have been widely used in previous studies, such as in Liang et al. (2018, 2019); Wang et al. (2019); Hyyppä et al. (2020a, c), and are classified into three complexity categories, i.e., ’easy’, ’medium’, and ’difficult’ based on, mainly tree density and amount of vegetation under the canopy in the plots. The third test area, a spruce forest in Paloheinä, Finland (60.26°N, 24.92°E) was used only to test the loop detection performance of VINS-Fusion.
The first plot in Evo has been classified as ’medium’ (Figure 3), and the second has been classified as ’difficult’ (Figure 4). The medium-difficulty plot had a tree density of 650 trees/ha and a mean DBH of 28 cm. The most common tree species in the plot were Norway spruce (Picea abies (L.) H.Karst.)(81.2 %) and Scots pine (Pinus sylvestris L.)(9.4 %). The difficult forest had a tree density of 2000 trees/ha and a mean DBH of 17 cm. The most common tree species in the plot were Norway spruce (64.4 %) and aspen (Populus tremula, L.) (18.3 %). Only trees with a DBH greater than 5 cm were included in these statistics. From now on, these two test areas are called ’Evo-medium’ and ’Evo-difficult’, respectively.
Figure 3: View from the first takeoff location in the ’Evo-medium’ test area. The takeoff location was changed after five test flights.
Figure 4: View from the takeoff location in the ’Evo-difficult’ test forest.
For reference purposes, circular GCPs provided by Agisoft Metashape (Agisoft LLC, St. Petersburg, Russia) were placed within the forest plots used in the flight tests. The targets had a diameter of 24.7 cm and the middle dots had a size of 7 cm. Relative distances between the GCPs were measured with Leica Nova TS60 total station with a precision of approximately 2 mm (Leica Geosystems 2020). Figure 5 shows GCPs in ’Evo-difficult’ during the test flights.
Figure 5: GCPs for reference trajectory generation in ’Evo-difficult’ forest. Targets were spread widely to the test site to increase the probability that there would be targets visible in the recorded camera data regardless the path that the robot takes.
GCPs were not used in the third test forest, ’Paloheinä’. The experiment there aimed to test the loop detection performance of VINS-Fusion in a homogeneous forest environment without any artificial targets. The data were collected by walking with the drone. The estimated average tree density in the area was approximately 2000 trees/ha.